# TrailBlazer: Trajectory Control for Diffusion-Based Video Generation

### 31 Dec 2023 v1

@@ Abstract
Within recent approaches to text-to-video (T2V) generation, achieving controllability in the synthesized video is often a challenge. Typically, this issue is addressed by providing low-level per-frame guidance in the form of edge maps, depth maps, or an existing video to be altered. However, the process of obtaining such guidance can be labor-intensive. This paper focuses on enhancing controllability in video synthesis by employing straightforward bounding boxes to guide the subject in various ways, all without the need for neural network training, finetuning, optimization at inference time, or the use of pre-existing videos. Our algorithm, TrailBlazer, is constructed upon a pre-trained (T2V) model, and easy to implement. [^1] Our project page: https://hohonu-vicml.github.io/Trailblazer.Page/ The subject is directed by a bounding box through the proposed spatial and temporal attention map editing. Moreover, we introduce the concept of keyframing, allowing the subject trajectory and overall appearance to be guided by both a moving bounding box and corresponding prompts, without the need to provide a detailed mask. The method is efficient, with negligible additional computation relative to the underlying pre-trained model. Despite the simplicity of the bounding box guidance, the resulting motion is surprisingly natural, with emergent effects including perspective and movement toward the virtual camera as the box size increases.
Figure 1: TrailBlazer extends a pre-trained video diffusion model to introduce trajectory control over one or multiple subjects. Its primary contribution lies in the ability to animate the synthesized subject using a bounding box (bbox), whether it remains static (Top-left) or dynamic in terms of location (Top-right), bbox size (Middle-left), and varied movement speed (Middle-right, the cat sitting in the early half of video in red bbox, and then moving with cyan bbox), achieved through keyframing. The moving subjects fit naturally within an environment specified by the overall prompt (Bottom-right). Additionally, the speed of the subjects can be controlled through keyframing (Bottom-left).
@@ 1 Introduction
Advancements in generative models for text-to-image (T2I) have been dramatic Ramesh et al. (2022); Saharia et al. (2022a); Rombach et al. (2022); Balaji et al. (2022). Recently, text-to-video (T2V) systems have made significant strides, enabling the automatic generation of videos based on textual prompt descriptions Ho et al. (2022a, b); Wu et al. (2023); Esser et al. (2023). One primary challenge in video synthesis lies in the extensive memory and training data required. Methods based on the pre-trained Stable Diffusion (SD) model have been proposed to address the efficiency issues in Text-to-Video (T2V) synthesis. These approaches address the problem from several perspectives including finetuning and zero-shot learning Khachatryan et al. (2023); Qi et al. (2023).
@@
However, text prompts do not provide good control over the spatial layout and trajectories of objects in the generated video. This control is known to be required for understandable narration of a story Arijon (1976). Existing work such as Hu and Xu (2023) has approached this problem by providing low-level control signals, e.g., using Canny edge maps or tracked skeletons to guide the objects in the video using ControlNet Zhang and Agrawala (2023). These methods achieve good controllability, but they can require considerable effort to produce the control signal. For example, capturing the desired motion of an animal (e.g., a tiger) or an expensive object (a jet plane) would be quite difficult, while sketching the desired movement on a frame-by-frame basis would be tedious.
@@
To address the needs of casual users, we introduce a high-level interface for control of object trajectories in synthesized videos. Users simply provide bounding boxes (bboxes) specifying the desired position of an object at several points in the video, together with the text prompt(s) describing the object at the corresponding times. Taking inspiration from the observation Liew et al. (2022) that object position is established early in the denoising process, and leveraging the clear spatial interpretation of spatial and temporal attention maps as illustrated in Fig. 2, we propose a general approach: Our strategy involves editing both spatial and temporal attention maps for a specific object during the initial denoising diffusion steps to concentrate activation at the desired object location. Our inference-time editing approach achieves this without disrupting the learned text-image association in the pre-trained model, and requires minimal code modifications. We use ZeroScope cerspense (2023) in a pre-trained fashion as the underlying model.
@@
  • Novelty. We introduce a novel approach employing high-level bounding boxes to guide the subject in diffusion-based video synthesis. This approach is suitable for casual users, as it avoids the need to record or draw a frame-by-frame positioning control signal. In contrast, the low-level guidance signals (detailed masks, edge maps) used by some other approaches have two disadvantages: it is difficult for non-artists to draw these shapes, and processing existing videos to obtain these signals limits the available motion to copies of existing sources.
  • Position, size, and prompt trajectory control. Our approach enables users to position the subject by keyframing its bounding box. The size of the bbox can be similarly controlled, thereby producing perspective effects (Figs. 1,6). Finally, users can also keyframe the text prompt to influence the behavior of the subject in the synthesized video (Figs. 1,6).
  • Simplicity. Our method operates by directly editing the spatial and temporal attention in the pre-trained denoising UNet. It requiring no training or optimization, and the core algorithm can be implemented in less than 200 lines of code.
@@ 2.1 Text-to-Image (T2I)
Denoising diffusion models construct a stochastic Sohl-Dickstein et al. (2015); Song and Ermon (2019); Ho et al. (2020) or deterministic Song et al. (2021) mapping between the data space and a corresponding-dimension multivariate Gaussian. Signals are synthesized by sampling from the Gaussian and performing a sequence of denoising steps. A number of works Nichol et al. (2022); Nichol and Dhariwal (2021); Ramesh et al. (2022); Saharia et al. (2022b) have performed T2I synthesis using images conditioned on the text embedding from a model such as CLIP Radford et al. (2021). Performance is significantly improved in the Latent Diffusion Model Rombach et al. (2022) (LDM) by doing the diffusion computation in the latent space of a carefully trained variational autoencoder. LDM was trained with a large scale dataset, resulting in the widely adopted Stable Diffusion (SD) system. We omit the basic diffusion derivation as good tutorials are available, e.g., Weng (2021).
Despite the success of image generation using SD, it is widely acknowledged that SD lacks controllability in synthesis. SD faces challenges in synthesizing multiple objects, often resulting in missing objects or incorrect assignment of prompt attributes to different objects. Recently, ControlNet Zhang and Agrawala (2023) and T2I-Adapter Mou et al. (2023) used additional fine-tuning layers to train the model with various forms of image conditioning (edge maps, skeletons).
The methods of Zhao et al. (2020); Sun and Wu (2022); Yang et al. (2022b); Ma et al. (2023); Xie et al. (2023); Bar-Tal et al. (2023) have addressed the layout-to-image (L2I) issue using few-shot learning. Directed Diffusion Ma et al. (2023), BoxDiff Xie et al. (2023), and MultiDiffusion Bar-Tal et al. (2023) use coarse bounding boxes to control subject position, achieving good result by manipulating the spatial latent and text embeddings cross attention map Hertz et al. (2022).
@@ 2.2 Text-to-Video (T2V)
Text-to-video (T2V) synthesis is generally more difficult than T2I due to the difficulty of ensuring temporal consistency and requirement for a large paired text and video dataset. Ho et al. (2022b); Harvey et al. (2022); Höppe et al. (2022); Voleti et al. (2022); Yang et al. (2022a); Ge et al. (2023) show methods that build on top of image diffusion models. Some works Blattmann et al. (2023); Luo et al. (2023) also introduce 3D convolutional layers in the denoising UNet to learn temporal information. Imagen Video Ho et al. (2022a) achieves higher resolution by computing temporal and spatial super-resolution on initial low resolution videos. VideoLDM Blattmann et al. (2023) and ModelScope Luo et al. (2023) insert a temporal attention layer by reshaping the latent tensor. Text2Video-Zero Khachatryan et al. (2023) and FateZero Qi et al. (2023) investigate how the temporal coherence can be improved by cross frame attention manipulation with pre-trained T2I models. Ge et al. (2023) address the same problem by introducing temporal correlation in the diffusion noise. However, these pioneering studies generally lack position control in the video synthesis.
Recently several works have been proposed to solve the controllability in video synthesis problem by using pre-trained models together with low-level conditioning information such as edge or depth maps. Control-A-Video Chen et al. (2023) and MagicProp Yan et al. (2023) use depth maps with ControlNet to train a temporal-aware network. Text2Video-Zero Khachatryan et al. (2023) partially achieves controllability by initializing the latent frames conditioned on the first frame with applied linear translation. However, it does not know about the reconstructed subject, making it difficult to edit the video (e.g., the user might need to know the initial position after synthesis, and then adjust the the shifting factor). Distinct from the methods above, we use an attention injection method to guide the denoising path rather than optimization, and in general this is robust to different random seeds. We note that our method can position and animate the bounding boxes with keyframing to control the subject.
@@ 3 Method
Our approach is based on the open-source pre-trained model ZeroScope cerspense (2023). This is a fine-tuned version of ModelScope Luo et al. (2023), known for its ability to generate high-quality videos without significant temporal flickering. It is noteworthy that our approach preserves this desirable temporal coherence effect achieved in their work. Our approach does not require any training, optimization, or low-level control signals (e.g., edge, depth maps with ControlNet Zhang and Agrawala (2023)). On the contrary, all that is required from the user is the prompt and an approximate bounding box (bbox) of the subject. Bboxes and corresponding prompts can be specified at several points in the video, and these are treated as keyframes and interpolated to smoothly control both the motion and prompt content. We use the following notation: Bold capital letters (e.g., $\mathbf{M}$) denote a matrix or a tensor depending on the context, vectors are represented with bold lowercase letters (e.g., $\mathbf{m}$), and scalars are denoted as lowercase letters (e.g., $m$). We use superscripts to denote an indexed tensor slice (e.g., $\mathbf{M}^{(i)}$). A synthesized video is composed of a number of images ordered in time. The individual images will be referred to as frames, and the collection of corresponding times is the timeline. Spatial or temporal attention will be informally referred to as correlation.
@@
Similar to the work in Ma et al. (2023), our method draws significant inspiration from visual inspection of cross-attention maps. Consider the final cross-attention result depicted in Fig. 2, generated from the prompt “an astronaut walking on the moon”. The spatial cross attention (denoted as SAttn) associated with the prompt word “astronaut” is highlighted in the second row, showcasing the overall position of the subject. Furthermore, we visualize the attention map from the temporal module in the pre-trained model. The third row displays “self-frame” temporal attention maps TAttn-Self, which consistently align with SAttn.
@@
The last row of Fig. 2 presents the visualization of cross-frame temporal attention maps, denoted as TAttn-Cross, illustrating the attention between the first frame and subsequent frames in the video. As the distance between frames increases, the attention becomes less correlated in the subject area and becomes more correlated in the background area. This observation aligns with the reconstructed video shown in the first row, where the background remains nearly static while the astronaut’s motion varies frame by frame. We will consider the temporal attention in detail in Sec. 3.3.
@@
Figure 2: Basis of our method. We draw inspiration from inspection of the spatial (SAttn) and temporal (TAttn) cross-attention maps viewed with self-frame attention (Self) and cross-frame attention (Cross).Thus TAttn-Self and TAttn-Cross denote the self- and cross-frame attention map, respectively. SAttn is the spatial cross-attention map with the prompt word “astronaut”. The symbol “Attn(i,j)” denotes the temporal attention map between frame $i$ and frame $j$. The first row shows reconstructions sampled from frames 1, 4, 16, and 24, respectively. In the TAttn-Cross images, the frame number were manually chosen to best illustrate the cross-frame attention between the astronaut and the background. Please refer to the main text for more details.
@@ 3.1 Pipeline
Figure 3: Pipeline Overview. Our pipeline highlights the central components of spatial cross-attention editing (left, in the blanched almond-colored section) and temporal cross-frame attention editing (right, in the blue section). This operation is exclusively applied during the denoising process in the early stage. The objective is to alter the attention map (e.g., $\mathbf{A}_{s},\mathbf{A}_{m}$) using a Gaussian weighting within a user-specified bbox.
As mentioned above, keyframing wiki (2023) is a technique that defines properties of images at particular frames (keys) in timeline and then automatically interpolates these values to achieve a smooth transition between the keys. It is widely used in the movie animation and visual effects industries since it reduces the artist’s work while simultaneously producing temporally smooth motion that would be hard to achieve if the artist directly edited every image. Our system takes the advantage of this principle, and asks the user to specify several keys, consisting of bboxes and the associated prompts, describing the subject location and appearance or behavior at the particular times. For instance, as shown in Fig. 1 (Middle-right), the video of the cat initially sitting on the left, then running to the right, is achieved simply by placing keys at three frames only. Specifically, the sitting cat in the first part of the video is obtained with two identically positioned bboxes on the left, with the keyframes at the beginning and middle of the timeline and the prompt word “sitting” associated with both. A third keyframe is placed at the end of the video, with the bbox positioned on the right together with the prompt changing to “running”. This results in the cat smoothly transitioning from sitting to running in the second part of the video.
We use the pre-trained ZeroScope model cerspense (2023) in all our experiments with no neural network training, finetuning, or optimization at inference time. Our pipeline is shown in Fig. 3. The spatial cross attention and the temporal attention is discussed in detail in the Sec. 3.2 and Sec. 3.3, respectively. The editing of all the spatial and the temporal editing is performed in the early steps $t\in\{T,...,T-N_{S}\}$, and $t\in\{T,...,T-N_{M}\}$ of backward denoising process, where $T$ is the total number of denoising time steps, and $N_{S}$, and $N_{M}$ are hyperparameters specifying the number of steps of spatial and temporal attention editing. The parameter setting is detailed in our supplementary material.
@@
Our system processes a set of keyframes, encompassing associated bbox regions $\mathcal{R}_{f}$ and prompts $\mathcal{P}_{f}$ at frame $f$, where $f$ denotes the frame index within the range $f\in\{0,...,N_{F}\}$. In practice, users are required to specify a minimum of two keyframes at the start and end of the video sequence. Then, the information in these keyframes is linearly interpolated, such as the bbox $\mathcal{B}_{f}$ and the prompt text embedding $y(\mathcal{P}_{f})$ through the text encoder $y(\cdot)$. To enhance readability, we omit subscript $f$ and the linearly blended video sequence between the keyframes when discussing the core method.
The region $\mathcal{R}$ is characterized of a set of parameters =$\mathcal{R}=\{\mathcal{B},\mathcal{I},\mathcal{T}\}$: a set of bbox positions (e.g., $\mathcal{B}$), the indices of the subject we would like to constrain (e.g., $\mathcal{I}$), and the indices of the trailing maps (e.g., $\mathcal{T}$) to enhance the controllability. The subject indices $\mathcal{I}\subset\{i|i\in\mathbb{N},1\leq i\leq|\mathcal{P}|\}$, are 1-indexed with the associated word in the prompt. For example, =$\mathcal{I}=\{1,2\}$ is associated with “a”, “cat” in the prompt “a cat sitting on the car”. The trailing attention indices $\mathcal{T}$ and the bounding box (bbox) $\mathcal{B}$ are defined as below.
The trailing attention maps indices $\mathcal{T}\subset\{i|i\in\mathbb{N},|\mathcal{P}|<i\leq N_{P}\}$ is the set of indices corresponding the cross-attention maps generated without a prompt word association, where $N_{P}$ denotes the maximum prompt length that a tokenizer model can take, which is $N_{P}=77$ when CLIP is used Radford et al. (2021). It serves as a means of controlling the spatial location of the synthesized subject and its attributes. A larger number of trailing indices set provides greater controllability but comes with the risk of failed reconstruction, as illustrated in Fig. 8. We refer reader to the work Ma et al. (2023) for more detail.
A bbox =$\mathcal{B}=\bigl{\{}(x,y)\,|\,b_{\text{left}}\times w\leq x\leq b_{\text{right}}\times w,\;b_{\text{top}}\times h\leq y\leq b_{\text{bottom}}\times h\bigl{\}}$, is a set of all pixel coordinates inside the bbox of resolution $w\times h$. In our implementation, $\mathcal{B}$ is produced by a tuple of the four scalars representing the boundary of the bbox $\mathbf{b}=(b_{\text{left}},b_{\text{top}},b_{\text{right}},b_{\text{bottom}})$, where $b_{\text{left}},b_{\text{top}},b_{\text{right}},b_{\text{bottom}}\in[0,1]$ in the ratio of the synthesis resolution. The height and the width, denoted as $h$ and $w$ respectively, are defined by the resolution of the UNet intermediate representation Rombach et al. (2022).

### 08 Apr 2024 v2

@@ Abstract
Within recent approaches to text-to-video (T2V) generation, achieving controllability in the synthesized video is often a challenge. Typically, this issue is addressed by providing low-level per-frame guidance in the form of edge maps, depth maps, or an existing video to be altered. However, the process of obtaining such guidance can be labor-intensive. This paper focuses on enhancing controllability in video synthesis by employing straightforward bounding boxes to guide the subject in various ways, all without the need for neural network training, finetuning, optimization at inference time, or the use of pre-existing videos. Our algorithm, TrailBlazer, is constructed upon a pre-trained (T2V) model, and easy to implement.  [^1] Our project page: https://hohonu-vicml.github.io/Trailblazer.Page/ The subject is directed by a bounding box through the proposed spatial and temporal attention map editing. Moreover, we introduce the concept of keyframing, allowing the subject trajectory, morphing, and overall appearance to be guided by both a moving bounding box and corresponding prompts, without the need to provide a detailed mask. The method is efficient, with negligible additional computation relative to the underlying pre-trained model. Despite the simplicity of the bounding box guidance, the resulting motion is surprisingly natural, with emergent effects including perspective and movement toward the virtual camera as the box size increases.
Figure 1: TrailBlazer extends a pre-trained video diffusion model to introduce trajectory control over one or multiple subjects. Its primary contribution lies in the ability to animate the synthesized subject using a bounding box (bbox), whether it remains static (Top-left) or dynamic in terms of location and bbox size (Top-right), morphing for subject interpolation (Middle-left), and varied movement speed (Middle-right). The moving subjects fit naturally within an environment specified by the overall prompt (Bottom-right). Additionally, the speed of the subjects can be controlled through keyframing (Bottom-left).
@@ 1 Introduction
Advancements in generative models for text-to-image (T2I) have been dramatic Ramesh et al. (2022); Saharia et al. (2022a); Rombach et al. (2022); Balaji et al. (2022). Recently, text-to-video (T2V) systems have made significant strides, enabling the automatic generation of videos based on textual prompt descriptions Ho et al. (2022a, b); Wu et al. (2023); Esser et al. (2023). One primary challenge in video synthesis lies in the extensive memory and required training data. Methods based on the pre-trained Stable Diffusion (SD) model have been proposed to address the efficiency issues in T2V synthesis. These approaches address the problem from several perspectives including finetuning and zero-shot learning Khachatryan et al. (2023); Qi et al. (2023).
@@
However, text prompts do not provide good control over the spatial layout and trajectories of objects in the generated video. This control is known to be required for understandable narration of a story Arijon (1976). Existing work such as Hu and Xu (2023) has approached this problem by providing low-level control signals, e.g., using Canny edge maps or tracked skeletons to guide the objects in the video using ControlNet Zhang and Agrawala (2023). These methods achieve good controllability, but they can require considerable effort to produce the control signal. For example, capturing the desired motion of an animal (e.g., a tiger) or an expensive object (e.g., a jet plane) would be quite difficult, while sketching the desired movement on a frame-by-frame basis would be tedious.
To address the needs of casual users, we introduce a high-level interface for the control of object trajectories in synthesized videos summarized in Fig. 1. Users simply provide bounding boxes (bboxes) specifying the desired position of an object at several times (keyframes) in the video, together with the text prompt(s) describing the object at the corresponding times. The provided bboxes are interpolated between the keyframes, resulting in smooth motion and size changes of the object. For instance, the cat sitting in the early half of the video in the red bbox, and then moving with the cyan bbox achieved through keyframing in the middle right of Fig. 1.
@@
If more than one different text prompt is provided, these prompts are also interpolated, resulting in a “morphing” effect such as the cat that transforms into a dog in Fig. 1. To achieve this guidance we take inspiration from the observation Liew et al. (2022) that object position is established early in the denoising diffusion process, and we leverage the clear spatial interpretation of spatial and temporal attention maps as illustrated in Fig. 2. Our resulting strategy involves editing both spatial and temporal attention maps for a specific object during the initial denoising diffusion steps to concentrate activation at the desired object location. Our inference-time editing approach achieves this without disrupting the learned text-image association in the pre-trained model, and requires minimal code modifications.
Our method, TrailBlazer, builds on previous works. We use the pre-trained ZeroScope model cerspense (2023), which is a fine-tuned version of Wang et al. (2023), as our underlying model. A body of previous and concurrent works have addressed guiding object position in image generation models, including Zhao et al. (2020); Sun and Wu (2022); Yang et al. (2022b); Balaji et al. (2022); Ma et al. (2023); Xie et al. (2023); Li et al. (2023); Bar-Tal et al. (2023). TrailBlazer most closely resembles the cross-attention injection used in Ma et al. (2023), and we adopt some notation from that paper. However, our work addresses a different problem, that of controlling position and trajectories in videos, which requires a different approach to control temporal cross-frame attention. Our work also does not require the inference-time optimization algorithm used in Ma et al. (2023).
@@
  • Novelty. We introduce a novel approach employing high-level bounding boxes to guide the subject in diffusion-based video synthesis. This approach is suitable for casual users, as it avoids the need to record or draw a frame-by-frame positioning control signal. In contrast, the low-level guidance signals such as detailed masks, edge maps, used by some other approaches have two disadvantages: it is difficult for non-artists to draw these shapes, and processing existing videos to obtain these signals limits the available motion to copies of existing sources.
  • Position, size, and prompt trajectory control. Our approach enables users to position the subject by keyframing its bounding box. The size of the bbox can be similarly controlled, thereby producing perspective effects (Figs. 1,6). Finally, users can also keyframe the text prompt to influence the behavior and identity of the subject in the synthesized video (Figs. 1).
  • Simplicity. Our method operates by directly editing the spatial and temporal attention in the pre-trained denoising UNet. It requires no training or optimization, and the core algorithm can be implemented in less than 200 lines of code.
@@ 2.1 Text-to-Image (T2I)
Denoising diffusion models construct a stochastic Sohl-Dickstein et al. (2015); Song and Ermon (2019); Ho et al. (2020) or deterministic Song et al. (2021) mapping between the data space and a corresponding-dimension multivariate Gaussian. Signals are synthesized by sampling from a normal distribution and performing a sequence of denoising steps. A number of works Nichol et al. (2022); Nichol and Dhariwal (2021); Ramesh et al. (2022); Saharia et al. (2022b) have performed T2I synthesis using images conditioned on the text embedding from a model such as CLIP Radford et al. (2021). Performance is significantly improved in the Latent Diffusion Model Rombach et al. (2022) (LDM) by performing the diffusion computation in the latent space of a carefully trained variational autoencoder. LDM was trained with a large scale dataset, resulting in the widely adopted Stable Diffusion (SD) system. We omit the basic diffusion derivation as tutorials are available, e.g., Weng (2021).
Despite the success of image generation using SD, it is widely acknowledged that SD lacks controllability in synthesis. SD faces challenges in synthesizing multiple objects, often resulting in missing objects or incorrect assignment of prompt attributes to different objects. Recently, ControlNet Zhang and Agrawala (2023) and T2I-Adapter Mou et al. (2023) introduced additional fine-tuning layers to train the model with various forms of image conditioning such as edge maps, or rigging skeletons.
The methods of Zhao et al. (2020); Sun and Wu (2022); Yang et al. (2022b); Ma et al. (2023); Xie et al. (2023); Bar-Tal et al. (2023) have addressed the layout-to-image (L2I) issue using few-shot learning. Directed Diffusion Ma et al. (2023), BoxDiff Xie et al. (2023), and MultiDiffusion Bar-Tal et al. (2023) use coarse bboxes to control subject position, achieving good results by manipulating the spatial latent and text embeddings cross attention map Hertz et al. (2022).
@@ 2.2 Text-to-Video (T2V)
Text-to-video (T2V) synthesis is generally more difficult than T2I due to the difficulty of ensuring temporal consistency and the requirement for a large paired text and video dataset. Ho et al. (2022b); Harvey et al. (2022); Höppe et al. (2022); Voleti et al. (2022); Yang et al. (2022a); Ge et al. (2023) show methods that build on top of image diffusion models. Some works Blattmann et al. (2023); Luo et al. (2023) also introduce 3D convolutional layers in the denoising UNet to learn temporal information. Imagen Video Ho et al. (2022a) achieves higher resolution by computing temporal and spatial super-resolution on initial low resolution videos. VideoLDM Blattmann et al. (2023) and ModelScope Luo et al. (2023) insert a temporal attention layer by reshaping the latent tensor. Text2Video-Zero Khachatryan et al. (2023), denoted as T2V-Zero, and FateZero Qi et al. (2023) investigate how the temporal coherence can be improved by cross frame attention manipulation with pre-trained T2I models. Ge et al. (2023) addresses the same problem by introducing temporal correlation in the diffusion noise. However, these pioneering studies generally lack position control in the video synthesis.
Recently several works have been proposed to solve the controllability in video synthesis problem by using pre-trained models together with low-level conditioning information such as edge or depth maps. Control-A-Video Chen et al. (2023) and MagicProp Yan et al. (2023) use depth maps with ControlNet to train a temporal-aware network. T2V-Zero Khachatryan et al. (2023) partially achieves controllability by initializing the latent frames conditioned on the first frame with applied linear translation. However, the control is indirect and requires two steps. First, the user first needs to locate the subject’s numerical location, and then adjust a translation offset. Distinct from the methods above, we use an attention injection method to guide the denoising path rather than optimization, and in general this is robust to different random seeds. The recent project Peekaboo Jain et al. (2023) is concurrent with TrailBlazer and shares similar goals. Both Peekaboo and TrailBlazer guide subjects in video by manipulating the attention, however the formulations differ in many details. Peekaboo’s use of an infinite negative attention injection in the background regions appears to often result in backgrounds with missing detail. In Sec. 4, we will provide both quantitative and visual evidence to demonstrate the better controllability and quality of our results. Other very recent preprints that address the layout-to-video problem in differing ways include Lian et al. (2023); Wang et al. (2024); Yang et al. (2024). We do not compare against these concurrent works because their source was not available at the time of writing.
@@ 3 Method
TrailBlazer is based on the open-source pre-trained model ZeroScope cerspense (2023). This is a fine-tuned version of ModelScope Luo et al. (2023), known for its ability to generate high-quality videos without significant temporal flickering. It is noteworthy that TrailBlazer preserves this desirable temporal coherence effect achieved in their work. TrailBlazer does not require any training, optimization, or low-level control signals (e.g., edge, depth maps with ControlNet Zhang and Agrawala (2023)). On the contrary, all that is required from the user is the prompt and an approximate bounding box (bbox) of the subject. Bboxes and corresponding prompts can be specified at several points in the video, and these are treated as keyframes and interpolated to smoothly control both the motion and prompt content.
We use the following notation: Bold capital letters (e.g., $\mathbf{M}$) denote a matrix or a tensor depending on the context, vectors are represented with bold lowercase letters (e.g., $\mathbf{m}$), and scalars are denoted as lowercase letters (e.g., $m$). We use superscripts to denote an indexed tensor slice (e.g., $\mathbf{M}^{(i)}$). A synthesized video is composed of a number of images ordered in time. The individual images will be referred to as frames, and the collection of corresponding times is the timeline. Spatial or temporal attention will be informally referred to as correlation.
@@
Similar to the work in Ma et al. (2023), our method draws significant inspiration from visual inspection of cross-attention maps. Consider the final cross-attention result depicted in Fig. 2, generated from the prompt “an astronaut walking on the moon”. The spatial cross attention, denoted as SA-Cross, associated with the prompt word “astronaut” is highlighted at the left of the second row, showcasing the overall position of the subject. Furthermore, we visualize the attention map from the temporal module in the pre-trained model. The right in the first row displays “self-frame” temporal attention maps, denoted as TA-Self, which consistently align with SA-Cross.
@@
The right in the second row of Fig. 2 presents the visualization of cross-frame temporal attention maps, denoted as TA-Cross, illustrating the attention between the first frame and subsequent frames in the video. As the distance between frames increases, the attention becomes less correlated in the subject area and becomes more correlated in the background area. This observation aligns with the reconstructed video shown in the left of the first row, where the background remains nearly static while the astronaut’s position varies frame by frame. We will consider the temporal attention in detail in Sec. 3.3.
@@
Figure 2: Basis of our method. We draw inspiration from inspection of the spatial (SA) and temporal (TA) attention maps viewed with self-frame attention (Self) and cross-frame attention (Cross). Thus, TA-Self and TA-Cross denote the self- and cross-frame attention map, respectively. SA-Cross is the spatial cross-attention map with the prompt word “astronaut”. The symbol “Attn(i,j)” denotes the temporal attention map between frame $i$ and frame $j$. The Recons subfigure shows reconstructions sampled from frames 1, 4, 16, and 24, respectively. In the TA-Cross, the frame number were manually chosen to best illustrate the cross-frame attention between the astronaut and the background. Please refer to the main text for more details.
@@ 3.1 Pipeline
Figure 3: Pipeline Overview. Our pipeline highlights the central components of spatial cross-attention editing (left, in the blanched almond-colored section) and temporal cross-frame attention editing (right, in the blue section). This operation is exclusively applied during the denoising process in the early stage. The objective is to alter the attention map (e.g., $\mathbf{A}_{s},\mathbf{A}_{m}$) using a Gaussian weighting within a user-specified bbox. This example uses one prompt word AttnMap and two trailing AttnMaps for guidance as highlighted in red.
As mentioned above, keyframing wiki (2023) is a technique that defines properties of images at particular frames (keys) in a timeline and then automatically interpolates these values to achieve a smooth transition between the keys. It is widely used in the movie animation and visual effects industries since it reduces the artist’s work while simultaneously producing temporally smooth motion that would be hard to achieve if the artist directly edited every image. Our system takes advantage of this principle, and asks the user to specify several keys, consisting of bboxes and the associated prompts, describing the subject location and appearance or behavior at the particular times. For instance, as shown in Fig. 1 (Middle-right), the video of the cat initially sitting on the left, then running to the right, is achieved simply by placing keys at three frames only. Specifically, the sitting cat in the first part of the video is obtained with two identically positioned bboxes on the left, with the keyframes at the beginning and middle of the timeline and the prompt word “sitting” associated with both. A third keyframe is placed at the end of the video, with the bbox positioned on the right together with the prompt changing to “running”. This results in the cat smoothly transitioning from sitting to running in the second part of the video.
We use the pre-trained ZeroScope model cerspense (2023) in all our experiments with no neural network training, finetuning, or optimization at inference time. Our pipeline is shown in Fig. 3. The spatial cross attention and the temporal attention is discussed in detail in the Sec. 3.2 and Sec. 3.3, respectively. All spatial and temporal editing is performed in the early steps $t\in\{T,...,T-N_{S}\}$, and $t\in\{T,...,T-N_{M}\}$ of the backward denoising process, where $T$ is the total number of denoising time steps, and $N_{S}$, and $N_{M}$ are hyperparameters specifying the number of steps of spatial and temporal attention editing. The parameter settings are detailed in our supplementary material.
@@
Our system processes a set of keyframes, encompassing associated bbox regions $\mathcal{R}_{f}$ and prompts $\mathcal{P}_{f}$ at frame $f$, where $f$ denotes the frame index within the range $f\in\{1,...,N_{F}\}$. The users are required to specify a minimum of two keyframes: one at the start and one at the end of the video sequence. The information in these keyframes is linearly interpolated, such as the bbox $\mathcal{B}_{f}$ and the prompt text embedding $y(\mathcal{P}_{f})$ through the text encoder $y(\cdot)$. To enhance readability, we omit the subscript $f$ and the linearly blended video sequence between the keyframes when discussing the core method.
A region $\mathcal{R}$ is characterized by a set of parameters =$\mathcal{R}=\{\mathcal{B},\mathcal{I},\mathcal{T}\}$: a set of bbox positions (e.g., $\mathcal{B}$), the indices of the subject we would like to constrain (e.g., $\mathcal{I}$), and the indices of the trailing maps (e.g., $\mathcal{T}$) to enhance the controllability. The subject indices $\mathcal{I}\subset\{i|i\in\mathbb{N},1\leq i\leq|\mathcal{P}|\}$, are 1-indexed with the associated word in the prompt. For example, =$\mathcal{I}=\{1,2\}$ is associated with “a”, “cat” in the prompt “a cat sitting on the car”.
The trailing attention maps indices $\mathcal{T}\subset\{i|i\in\mathbb{N},|\mathcal{P}|<i\leq N_{P}\}$ is the set of indices corresponding to the cross-attention maps generated without a prompt word association, where $N_{P}$ denotes the maximum prompt length that a tokenizer model can take, which is $N_{P}=77$ when CLIP is used Radford et al. (2021). The trailing attention maps serve as a means of controlling the spatial location of the synthesized subject and its attributes. A larger trailing indices set $|\mathcal{T}|$ provides greater controllability but comes with the risk of failed reconstruction Ma et al. (2023).
A bbox =$\mathcal{B}=\bigl{\{}(x,y)\,|\,b_{\text{left}}\times w\leq x\leq b_{\text{right}}\times w,\;b_{\text{top}}\times h\leq y\leq b_{\text{bottom}}\times h\bigl{\}}$, is a set of all pixel coordinates inside the bbox of resolution $w\times h$. In our implementation, $\mathcal{B}$ is produced by a tuple of the four scalars representing the boundary of the bbox $\mathbf{b}=(b_{\text{left}},b_{\text{top}},b_{\text{right}},b_{\text{bottom}})$, where $b_{\text{left}},b_{\text{top}},b_{\text{right}},b_{\text{bottom}}\in[0,1]$ specify the bbox relative to the synthesis resolution. The height $h$ and width $w$, are defined by the resolution of the UNet intermediate representation Rombach et al. (2022).